suppose tan A =(3/4) and 180
Answer by stanbon(75887)

suppose tan A =(3/4) and 180 -----------
tan = y/x
In QIII, y and x are both negative.
y = -3 ; x = -4
r = sqrt[3^2+4^2] = 5
---------
Note:: cos(A) = x/r = -4/5
---------------
sin(1/2 A) = sqrt[(1-(-4/5))/2]
------------------
cos(1/2 A) = sqrt[1+(-4/5))/2]
-------------------
tan(1/2 A) = (1-cos(x))/sin(x) = (1-(-4/5))/(-3/5)
-----
Cheers,
Stan H.
------------